OPTICAL STARK SPECTROSCOPY OF CHLORO-METHYLENE, HCCl

XIUJUAN ZHUANG AND TIMOTHY C. STEIMLE, Department of Chemistry and Biochemistry, Arizona State University, Tempe, AZ 85287; \underline{ZHONG WANG}, Math and Sciences Department, Suffolk County Community College,East Campus, Riverhead, NY, 11901.

The optical spectrum of chloro-methylene, HCCl, has been studied for more than 40 years by both conventional^{a} and laser-based^{b}^{c}^{d} spectroscopy. Surprisingly, numerous visible bands have yet to be characterized, due in part to known perturbations. Furthermore, the permanent electric dipole moment, μ_{el}, for any state has yet to be determined. Here we report on the field-free and optical Stark spectrum of the Ã^{1}A′′(060)−X̃^{1}A′(000) band system. A cold molecular beam sample was produced by skimming the output of a pulsed discharge source and the spectrum recorded at a resolution of approximately 30 MHz via LIF detection. The field-free spectrum was analyzed to produce an improved set of spectroscopic parameters for the Ã^{1}A′′(060)state. The Stark induced shifts were analyzed to determine the values of the a-component of μ_{el} for the X̃^{1}A′(000)state of 0.498(8)D. Small perturbations in the Ã^{1}A′′(060)state will be described.

aA. J. Merer and D.N. Travis Can. J. Phys., 44 525 1966.

bM.Kakimoto, S.Saito and E. Hirota J.Mol.Spectrosc., 97 194 1983.

cB.-C.Chang and T. Sears J.Mol.Spectrosc., 173 391 1995.

dH. Fan, I. Ionescu, C. Annesley, J. Cummins, M. Bowers and S. A. Reid J.Mol.Spectrosc., 225 43 2004.